40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2190     Entry time: Thu May 17 10:20:04 2018
Author: awade 
Type: DailyProgress 
Category: Laser 
Subject: Power drop south laser 

There has been a distinct drop in laser power in the south path since yesterday.

It looks like the drop in power happened between GPS* 1209950712 s and 1209966050 s (18:24:54 and 22:40:32).  This was in the lab when nobody was there.

I've been tuning the temperature on the cavities to bring the beatnote into range and the current south laser slow voltage is -6.2759 V when the cavity is locked -- as an aside cavity heaters are set to 0.455 W diff heating and 0.56831 W common heating.   This is about where it was before and tuning the laser slow temperature around doesn't seem to get back to the original power levels as reported by the reflection and transmission PDs.  So it seems we are not in a mode hopping region.  I've attached a dataview screen shot showing that the prior to the drop the south trans PD was giving 4.25 V and the reflected 1.094 V (when locked).  After the change in power (after some unlock time for the cavities) the south trans PD was 3.11 V  and south refl PD was 0.488 V (when locked).

The points at which the cavity was unlocked show reflected power dropped from 3.81 V to 2.56 V.  This doesn't look great and its not clear what is going on with the laser.

For reference the current power going into each cavity is 3.1 mW for the north and 2.02 mW for the south.  Visibility is good on the south at 62% (Refl Vmin=0.592, Vmax =2.55  V)

Not clear what happen here, there is no mode cleaner or ISS applied in the south at the moment so seems like the only source of a power change would be a misalignment or a change in the laser.  Nobody has been in the lab since last night and looking around with the IR viewer there is no apparent clipping.

 

For now I think 3.1 mW for the north is a little bit too high.  I've turned down north power to 2 mW to match the south and we will watch the changes in power over the next few days in both lasers.

 

* Note: the GPS time on fb4 has now drifted by 9 days, so not clear if this is correct time, it is certainly in the early hours of this morning.

Attachment 1: Screenshot_from_2018-05-17_10-19-21.png  35 kB  Uploaded Thu May 17 11:30:41 2018  | Hide | Hide all
Screenshot_from_2018-05-17_10-19-21.png
ELOG V3.1.3-